2x+5x+4=25
16=4w+2w-2
31+4t-t=40
7a+16-3a=-4
7m+3-4m=-9

Answer by Alan3354(69443)

2x+5x+4=25
16=4w+2w-2
31+4t-t=40
7a+16-3a=-4
7m+3-4m=-9
------------
Is that all one problem?
If so, --> no solution.
--------
If they're separate problems, show your efforts.
email via the Thank You not for help, or to check your work.